# Disentangling quantum autoencoder

### v1

@@ VIII.4 Ising ansatz
Next, we study models with polynomial Lie algebra, which can generalize with only polynomial resource cost [28]. In particular, in Fig. 9 we show the training and test loss for DQAE applied to evolution with the Ising model $U_{\text{ising}}$ as shown in in Fig. 5c. Remarkable, for $N\geq 6$, we require only
@@
Figure 9: DQAE for disentangling Ising model evolution $U_{\text{ising}}$ as shown in Fig. 5c. a) Training and b) test loss for different qubit number $N$ as function of number of circuit parameters $M$, normalized by critical number of circuit parameters $M_{\text{c}}$. We have $L=2$ training states for $N=4$, and remarkably $L=1$ for $N=6,8$.
@@ VIII.5 Stabilizer states
Finally, we apply the DQAE to stabilizer states. We generate training and test data as detailed in Sec. VIII.1, where we choose $V_{\text{Cliff}}$ to be a random Clifford circuit, and the input single-qubit states as one of the $6$ single-qubit stabilizer states, i.e. the eigenstates of the $\sigma^{x}$, $\sigma^{y}$ and $\sigma^{z}$ Pauli operators. The DQAE is trained with Algorithm 1. We show the test error for the DQAE after learning stabilizer states in Fig. 10. We find that the training with the algorithm nearly always converges, and the test error is only limited by the number of training states $L$. We numerically find that for at least up to $N\leq 10$ qubits the test errors collapse to a single curve when $C_{\text{test}}$ is normalized by the test error without training $C_{\text{test}}^{0}$ and we shift $L$ by $N/4$. This implies that the number of training states needed to generalize scales as
@@
Notably, the DQAE trained on stabilizer states is also able to disentangle non-stabilizer states of the form $V_{\text{Cliff}}\bigotimes_{j}|\phi_{j}\rangle$, where $|\phi_{j}\rangle$ are arbitrary (non-stabilizer) single-qubit states. This highlights that the DQAE is capable of out-of-distribution generalization [32, 28], i.e. it can generalize even when test and training data are drawn from different distributions.
@@
Figure 10: Test error $C_{\text{test}}$ for DQAE for stabilizer states as a function of number of training states $L$. DQAE is trained using Metropolis algorithm 1. We plot $C_{\text{test}}$ normalized to test error without training $C_{\text{test}}^{0}$.
@@ IX Discussion
We introduced the DQAE to disentangle quantum states into single-qubit product states. Entangled states are highly susceptible to noise due to the collapse of the wavefunction, while product stares are more robust. This allows the DQAE to significantly enhance noise robustness of quantum information processing, such as the transfer of quantum information [51]. In particular, for the transfer of entangled states via a qubit loss channel, the DQAE gives an exponential advantage in the number of copies compared to unencoded transfer.
@@
The DQAE can be trained with a simple cost function. We characterize overparameterization and generalization using the DQFIM, finding that for product states evolved with an Ising-type evolution, $M=O(N)$ circuit parameters and $L=1$ training data is sufficient for generalization. We also study DQAE for stabilizer states, finding that $L=O(N)$ training data are sufficient to generalize. Notably, we also showcase out-of-distribution generalization [32, 28], demonstrating that training on stabilizer states allows the DQAE to also disentangle a class of non-stabilizer states. Our algorithm for disentangling stabilizer states so far rests on a Metropolis type algorithm. We believe that a direct construction of the disentangling Clifford unitary can be found, similar to other dynamical learning problems for Cliffords [47, 52, 53].
@@
Although we considered the DQAE for a qubit-loss model or (unrecoverable) leakage errors, it can also be applied to quantum error detection [2, 54]: In this case, $N$ logical qubits are redundantly encoded into many physical qubits. By performing check measurements, one can detect whether a logical error has occurred, and which qubit has been affected. Let us assume that each logical qubit has probability $p_{\text{L}}$ of experiencing an uncorrectable logical error. Without DQAE, one has to discard the full logical state if one logical qubit fails. In contrast, for states disentangled by DQAE, only the affected logical qubit has to be discarded, while the other logical qubits are unaffected. Thus, the DQAE could help to reduce the effective error rate for fault-tolerant quantum information transfer and storing information in quantum memory.
@@
We note that the disentangling of quantum states could not only be used for noise suppression, but also for learning compact representations of quantum states [55]. Here, the DQAE can serve as a routine in quantum machine learning algorithms to learn simple representation of states.
@@
Alternatively, the DQAE could be used to enhance cross-platform verification protocols [56]. In particular, measuring the fidelity $|\langle\psi_{1}|\psi_{2}\rangle|^{2}$ between two $N$-qubit quantum states $|\psi_{1}\rangle$, $|\psi_{2}\rangle$ on two different devices (without a quantum channel between them) requires exponential complexity $O(2^{N})$ in general [57]. Instead, one can first apply the DQAE $U$ on both states to disentangle the states into single-qubit product states $U|\psi_{1}\rangle$, $U|\psi_{2}\rangle$. Note that this leaves the fidelity invariant. Then, we can efficiently verify the fidelity between the states by local tomography of the $N$ single-qubit states which scales only as $O(N)$.
@@
Finally, we comment on the relationship of DQAE to the recently proposed Clifford-enhanced matrix product state (CEMPS) [58] and their disentangling algorithms [58, 59, 60, 61]. Matrix product states (MPSs) are classical state representations with low entanglement, which can be combined with highly entangling Cliffords to be more expressible [58]. Clifford-disentangling of MPS is a classical simulation method that disentangles a (single) quantum state into a CEMPS. Such methods have applications for finding ground states [62] and simulating dynamics [63, 64, 65]. Notably, Clifford circuits doped with up to $N$ T-gates can be disentangled into a single-qubit product state with a Clifford unitary applied to it [59, 60]. We note that in contrast to Clifford-disentangling of MPS, DQAE is targeted to be trainable on quantum computers and can disentangle a large class of input states at the same time.
@@
Future work could implement our DQAE in experiment. The DQAE can be straightforwardly implemented, following experimental demonstrations of other types of quantum autoencoders [18, 17, 15, 66].
@@ Appendix A Increasing model capacity
In our discussion of the DQAE, we so far focused on mapping a dataset of $N$-qubit quantum states to $N$-qubit single-qubit product states. We can generalize this concept to involve a mapping from $N+K$ qubit states to an $N$-qubit product states, where $K\geq 0$. The additional $K$ qubits are mapped to a fixed $K$-qubit state $|a\rangle$, referred to as the reference state.
@@
We propose an alternative approach, where instead of mapping the last $K$ qubits of every state in our dataset to a fixed reference state $|a\rangle$, we allow it to occupy a set of orthonormal states, such as the $K$-qubit computational basis states $\{|i\rangle\}_{i=1}^{2^{K}}$. As we will see, this enhances the capacity of the autoencoder. After measurement of the $K$ qubits, the orignal state can be reconstructed by a modified entangler: One prepares $K$ additional qubits which are initialized with the measurement outcomes $c_{k}$ of the disentangler. Refer to Fig. 11 for a sketch of the capacity-enhanced DQAE.
@@
Figure 11: Illustration of an increased capacity DQAE that encodes a $N+K$-qubit state $|\psi\rangle$ into a product state comprising a $N$-qubit single-qubit product quantum state and $K$ classical bits $c_{k}=\{0,1\}$ by measuring the first $K$ qubits in the computational basis. The entangler $U^{\dagger}(\theta)$ takes as input computational basis states initialized with the $K$ measurement outcomes, along with the $N$ single-qubit product states, to reconstruct the original state $|\psi\rangle$.

### v2

@@ VIII.4 Ising ansatz
Next, we study models with polynomial Lie algebra, which can generalize with only polynomial resource cost [28]. In particular, in Fig. 9 we show the training and test loss for DQAE applied to evolution with the Ising model $U_{\text{ising}}$ as shown in Fig. 5c. Remarkably, for $N\geq 6$, we require only
@@
Figure 9: DQAE for disentangling Ising model evolution $U_{\text{ising}}$ as shown in Fig. 5c. a) Training and b) test loss for different qubit number NN as function of number of circuit parameters $M$, normalized by critical number of circuit parameters $M_{\text{c}}$. We have $L=2$ training states for $N=4$, and remarkably $L=1$ for $N=6,8$.
@@ VIII.5 Stabilizer states
Finally, we apply the DQAE to stabilizer states. We generate training and test data as detailed in Sec. VIII.1, where we choose $V_{\text{Cliff}}$ to be a random Clifford circuit, and the input single-qubit states as one of the $6$ single-qubit stabilizer states, i.e. the eigenstates of the $\sigma^{x}$, $\sigma^{y}$, and $\sigma^{z}$ Pauli operators. The DQAE is trained with Algorithm 2. We show the test error for the DQAE after learning stabilizer states in Fig. 10. We find that the training with the algorithm nearly always converges, and the test error is only limited by the number of training states $L$. We numerically find that for at least up to $N\leq 10$ qubits, the test errors collapse to a single curve when $C_{\text{test}}$ is normalized by the test error without training $C_{\text{test}}^{0}$ and we shift $L$ by $N/4$. This implies that the number of training states needed to generalize scales as
@@
Notably, after successfully training the DQAE on stabilizer states, it is also able to disentangle non-stabilizer states of the form $V_{\text{Cliff}}\bigotimes_{j}|\phi_{j}\rangle$, where $|\phi_{j}\rangle$ are arbitrary (non-stabilizer) single-qubit states. The fact that for this case generalization can extend beyond stabilizer states has been proven in Ref. [32]. Thus, DQAE is capable of out-of-distribution generalization [32, 28], i.e. it can generalize even when test and training data are drawn from different distributions.
@@
Figure 10: Test error $C_{\text{test}}$ for DQAE for stabilizer states as a function of the number of training states $L$. DQAE is trained using the Metropolis algorithm 2. We plot $C_{\text{test}}$ normalized to test error without training $C_{\text{test}}^{0}$.
@@ IX Discussion
We introduced the DQAE to disentangle quantum states into single-qubit product states. Entangled states are highly susceptible to noise due to the collapse of the wavefunction, while product states are more robust. This allows the DQAE to significantly enhance noise robustness of quantum information processing, such as the transfer of quantum information [51]. In particular, for the transfer of entangled states via a qubit loss channel, the DQAE gives an exponential advantage in the number of copies compared to unencoded transfer.
@@
The DQAE can be trained with a simple cost function. We characterize overparameterization and generalization using the DQFIM, finding that for product states evolved with an Ising-type evolution, $M=O(N)$ circuit parameters and $L=1$ training data is sufficient for generalization. We also study DQAE for stabilizer states, finding that $L=O(N)$ training data are sufficient to generalize. Notably, we also showcase out-of-distribution generalization [32, 28], demonstrating that training on stabilizer states allows the DQAE to also disentangle a class of non-stabilizer states. Our algorithm for disentangling stabilizer states so far rests on a Metropolis-type algorithm. We believe that a direct construction of the disentangling Clifford unitary can be found, similar to other dynamical learning problems for Cliffords [47, 52, 53].
@@
Although we considered the DQAE for a qubit-loss model or (unrecoverable) leakage errors, it can also be applied to quantum error detection [2, 54]: In this case, $N$ logical qubits are redundantly encoded into many physical qubits. By performing check measurements, one can detect whether a logical error has occurred and which qubit has been affected. Let us assume that each logical qubit has a probability $p_{\text{L}}$ of experiencing an uncorrectable logical error. Without DQAE, one has to discard the full logical state if one logical qubit fails. In contrast, for states disentangled by DQAE, only the affected logical qubit has to be discarded, while the other logical qubits are unaffected. Thus, the DQAE could help to reduce the effective error rate for fault-tolerant quantum information transfer and storing information in quantum memory. Of course, running the DQAE circuit itself can induce noise, especially on NISQ devices, which have imperfect gates. For the DQAE to effectively reduce noise, we must assume that the (noisy) DQAE circuit itself induces fewer errors than the loss channel. Thus, in practice, one has to consider this tradeoff by taking the actual noisy gate implementation into account.
So far, we assumed that qubit loss channels act identically across all qubits. In practice, however, loss rates often vary across different qubits, with loss $q_{j}$ for the $j$th qubit. Our results still apply when the qubit loss varies: In this case, Eq. (11) is an upper bound on $R$ where one chooses $q=\max_{j}q_{j}$.
@@
We note that the disentangling of quantum states could not only be used for noise suppression, but also for learning compact representations of quantum states [55]. Here, the DQAE can serve as a routine in quantum machine learning algorithms to learn a simpler representation of states.
@@
Alternatively, the DQAE could be used to enhance cross-platform verification protocols [56]. In particular, measuring the fidelity $|\langle\psi_{1}|\psi_{2}\rangle|^{2}$ between two $N$-qubit quantum states $|\psi_{1}\rangle$, $|\psi_{2}\rangle$ on two different devices (without a quantum channel between them) requires exponential complexity $O(2^{N})$ in general [57]. Instead, one can first apply the DQAE $U$ on both states to disentangle the states into single-qubit product states $U|\psi_{1}\rangle$, $U|\psi_{2}\rangle$. Note that this leaves the fidelity invariant. Then, we can efficiently verify the fidelity between the states by local tomography of the $N$ single-qubit states, which scales only as $O(N)$.
@@
Finally, we comment on the relationship of DQAE to the recently proposed Clifford-enhanced matrix product state (CEMPS) [58] and their disentangling algorithms [58, 59, 60, 61, 62]. Matrix product states (MPSs) are classical state representations with low entanglement, which can be combined with highly entangling Cliffords to be more expressible [58]. Clifford-disentangling of MPS is a classical simulation method that disentangles a (single) quantum state into a CEMPS. Such methods have applications for finding ground states [63] and simulating dynamics [64, 65, 66]. Notably, Clifford circuits doped with up to $N$ T-gates can be disentangled into a single-qubit product state with a Clifford unitary applied to it [59, 60, 61]. We note that, in contrast to Clifford-disentangling of MPS, DQAE is targeted to be trainable on quantum computers and also can disentangle a large class of different input states at the same time.
@@
Future work could implement our DQAE in an experiment on quantum hardware. The DQAE can be straightforwardly implemented, following experimental demonstrations of other types of quantum autoencoders [18, 17, 15, 67].
Data availability statement
The data cannot be made publicly available upon publication because they are not available in a format that is sufficiently accessible or reusable by other researchers. The data that support the findings of this study are available upon reasonable request from the authors.
@@ Appendix A Increasing model capacity
In our discussion of the DQAE, we have so far focused on mapping a dataset of $N$-qubit quantum states to $N$-qubit single-qubit product states. We can generalize this concept to involve a mapping from $N+K$ qubit states to $N$-qubit product states, where $K\geq 0$. The additional $K$ qubits are mapped to a fixed $K$-qubit state $|a\rangle$, referred to as the reference state.
@@
We propose an alternative approach, where instead of mapping the last $K$ qubits of every state in our dataset to a fixed reference state $|a\rangle$, we allow it to occupy a set of orthonormal states, such as the $K$-qubit computational basis states $\{|i\rangle\}_{i=1}^{2^{K}}$. As we will see, this enhances the capacity of the autoencoder. After measurement of the $K$ qubits, the original state can be reconstructed by a modified entangler: One prepares $K$ additional qubits which are initialized with the measurement outcomes $c_{k}$ of the disentangler. Refer to Fig. 11 for a sketch of the capacity-enhanced DQAE.
@@
Figure 11: Illustration of an increased capacity DQAE that encodes a $N+K$-qubit state $|\psi\rangle$ into a product state comprising a NN-qubit single-qubit product quantum state and $K$ classical bits $c_{k}=\{0,1\}$ by measuring the first $K$ qubits in the computational basis. The entangler $U^{\dagger}(\theta)$ takes as input computational basis states initialized with the $K$ measurement outcomes, along with the $N$ single-qubit product states, to reconstruct the original state $|\psi\rangle$.